Things did not go quite as planned. You invested $20,000, part of it in a stock that paid 12% annual interest. However, the rest of the money suffered a 5% loss. If the total annual income from both investments was $2026, how much was invested at each rate?

Found 2 solutions by Theo, solver91311:
Answer by Theo(13342)

total amount of money that was invested was 20,000.

one part of it earned 12%.

the other part lost 5%.

x = the part that earned 12%.
y = the part that lost 5%.

the total income from both investments was 2026.

your equation is:

.12 * x - .05 * y = 2026

also, x + y = 20,000

from the second equation, solve for y to get y = 20,000 - x

in the first equation replace y with 20,000 - x to get:

.12 * x - .05 * (20,000 - x) = 2026

simplify to get:

.12 * x - .05 * 20,000 + .05 * x = 2026

combine like terms to get:

.17 * x - 1000 = 2026

add 1000 to both sides of this equation to get:

.17 * x = 3026

solve for x to get x = 3026 / .17 = 17,800.

since x + y = 20,000, and x = 17,800, then y must be equal to 2,200.

he earned 12% on 17,800 and lost 5% on 2,200.

his net return was 2136 - 110 = 2026

that agrees with the problem statement, so the solution looks good.

the solution is he invested 17,800 that earned 12% and 2,200 that lost 5%.





Answer by solver91311(24713)





Solve for then calculate


John

My calculator said it, I believe it, that settles it